Please help me solve this problem, the problem states to solve: 6x^2+13x-5=0
Answer by stanbon(75887)

solve: 6x^2+13x-5=0
6x^2+15x-2x-5 = 0
3x(2x+5)-(2x+5)=0
(2x+5)(3x-1)=0
x = -5/2 or x=1/3
==================
Cheers,
Stan H.